fee simple interest
$300 is borrowed the amount owed after 12 months is $00 what is the interest rate.
25 is 12% of what number?
X + Y =130
6.50X +250Y - 604.5
Graph the solutions of equalities
4X+5Y<20
4x-2y<8

Answer by drk(1908)

You have several questions here. Let's take one at a time.
QUESTION #1:
$300 is borrowed the amount owed after 12 months is $00 what is the interest rate.
We need I = PRT for this question.
P = 300, t = 12 months. Time is measured in years, so t = 1. We don't have interest rate or amount of interest paid back. It appears according to the question that $300 is paid back after 12 months. Without some R or I info. I can't solve this one.
--
QUESTION #2: 25 is 12% of what number?
rewrite as
25 = (12/100)*N
cross multiply to get
2500 = 12N
divide to get
N = 208.33
--
QUESTION #3:
X + Y =130
6.50X +250Y - 604.5
Don't know if we are solving algebraically or graphing. I will show solve algebraically. I think you mean
(i) X + Y =130
(ii) 6.50X +2.50Y = 604.5
step 1: multiply (i) by -2.50 to get
-2.5X - 2.5Y = -325
6.50X + 2.50Y = 604.5
step 2: add down to eliminate Y.
4X = 279.5
step 3: solve for X.
X = 69.875.
step 4: solve for Y.
Y = 60.125
--
QUESTION #4:
Graph the solutions of equalities
(i) 4X + 5Y < 20
(ii) 4X - 2Y < 8
Find x and y intercepts for both equations.
(i) X-int = 5 ; Y-int = 4. Plot 5 on the X-axis and 4 on the Y-axis and connect with dashed line. Shade down toward you.
(ii) X-int = 2 ; Y-int = -4. Plot 2 on the X-axis and -4 on the Y-axis and connect with dashed line. Shade up away from you.
where there is double shading, you have the answer.